# SFO: Piloting VLM Feedback for Offline RL

### v2

@@ Abstract
While internet-scale image and textual data have enabled strong generalization in Vision-Language Models (VLMs), the absence of internet-scale control data has impeded the development of similar generalization in standard reinforcement learning (RL) agents. Although VLMs are fundamentally limited in their ability to solve control tasks due to their lack of action-conditioned training data, their capacity for image understanding allows them to provide valuable feedback in RL tasks by recognizing successful outcomes. A key challenge in Reinforcement Learning from AI Feedback (RLAIF) is determining how best to integrate VLM-derived signals into the learning process. We explore this question in the context of offline RL and introduce a class of methods called Sub-Trajectory Filtered Optimization (SFO). We identify three key insights. First, trajectory length plays a crucial role in offline RL, as full-trajectory preference learning exacerbates the stitching problem, necessitating the use of sub-trajectories. Second, even in Markovian environments, a non-Markovian reward signal from a sequence of images is required to assess trajectory improvement, as VLMs do not interpret control actions and must rely on visual cues over time. Third, a simple yet effective approach—filtered and weighted behavior cloning—consistently outperforms more complex RLHF-based methods. We propose Sub-Trajectory Filtered Behavior Cloning (SFBC), a method that leverages VLM feedback on sub-trajectories while incorporating a retrospective filtering mechanism that removes sub-trajectories preceding failures to improve robustness and prevent turbulence. This study is preliminary; we provide initial evidence through evaluations on a toy control domain. Please enjoy our airport puns.
@@ 1 Introduction
Reinforcement Learning from AI Feedback (RLAIF) presents a promising approach by leveraging vision-language models (VLMs) as scalable sources of learning signals (Bai et al.,, 2022; Klissarov et al.,, 2023; Ma et al.,, 2023; Faldor et al.,, 2024; Rocamonde et al.,, 2024; Baumli et al.,, 2023; Lee et al.,, 2024; Wang et al.,, 2024). However, determining how best to integrate VLM-derived signals into RL remains an open challenge. A key difficulty is that VLMs struggle to discriminate between similar trajectories, making it difficult to provide precise, informative feedback. This challenge is further complicated in online RL, where agents must learn from randomly initialized policies, leading to highly ambiguous early-stage feedback.
@@
    First, trajectory length is a crucial factor in offline RL—full-trajectory preference learning exacerbates the well-known stitching problem, necessitating the use of sub-trajectories, as depicted in Figure 1.
@@
    Second, even in Markovian environments, a non-Markovian reward model based on image sequences is required to assess trajectory improvement. Since VLMs are not trained on action-conditioned data, they have no basis for reasoning about how actions influence state transitions and must instead rely on visual cues over time. This is depicted in Figure 1. While discrete actions can sometimes be represented textually, many control tasks involve continuous actions (e.g., torques) that cannot be meaningfully encoded in a format suitable for VLMs. Furthermore, providing a reward at every time step by querying a VLM is computationally intractable for many problems without additional distillation Wang et al., (2024).
@@
To that end, we introduce Sub-Trajectory Filtered Behavior Cloning (SFBC), which leverages VLM feedback on sub-trajectories while incorporating a retrospective filtering mechanism that removes sub-trajectories preceding failures. We empirically evaluate SFBC in a toy control domain, demonstrating its effectiveness compared to existing baselines, including naive behavior cloning and preference-based methods such as Direct Preference Optimization (DPO) (Rafailov et al.,, 2023). Notably, our method outperforms AWAC (Nair et al.,, 2020) and TD3+BC (Fujimoto and Gu,, 2021) when using ground-truth rewards, highlighting the effectiveness of VLM-driven sub-trajectory feedback.
@@ 2 Related Work
One approach to improving RL efficiency is the manual collection of large-scale offline datasets (Vuong et al.,, 2023). However, this is difficult to scale due to the expense of generating diverse and high-quality trajectories on robotics in the real world. Similarly, human annotation of RL datasets is costly and difficult to scale, given the need for expert-level feedback and the challenge of obtaining consistent reward signals across tasks.
VLMs offer a promising alternative to human annotation by providing AI-generated feedback. Prior work has explored using VLMs for reward design and environment generation in code. Eureka and OMNI-EPIC (Ma et al.,, 2023; Faldor et al.,, 2024) have proposed methods where VLMs write programs to specify rewards and structure environments for learning. However, just as machine learning is needed to specify object categories (e.g., recognizing a dishwasher) rather than relying on hand-crafted rules, RL also requires machine-learned reward models rather than fixed rules from a VLM (e.g., to identify when a dishwasher has been unloaded). We firmly believe that hardcoded rules from a VLM cannot specify sufficiently complex tasks to elicit sufficiently robust reinforcement learning agents.
Recent work has also attempted to leverage AI-based reward detection to improve RLHF. Agent Q (Putta et al.,, 2024) uses AI-driven reward and success detection to generate preference labels for DPO. RL-VLM-F (Wang et al.,, 2024) distills a reward model from AI-generated labels but operates on Markovian states, which we find to be suboptimal. Rocamonde et al., (2024) and Baumli et al., (2023) use a CLIP model to provide Markovian feedback. Additionally, RoboCLIP (Sontakke et al.,, 2023) encodes text and video representations using CLIP to provide feedback. While RoboCLIP-style methods offer a useful way to encode feedback, we propose an end-to-end approach using a VLM (GPT-4o in our case), which provides greater flexibility. Our study focuses on how best to utilize this feedback and how to leverage feedback in an offline context.
Complementary approaches involve automatic environment generation and world modeling (Bruce et al.,, 2024; Valevski et al.,, 2024). These efforts provide a scalable way to generate diverse training environments, but they still require specifying reward signals to learn effectively. Similarly, automated task selection (Wang et al.,, 2023) and task ordering techniques such as Prioritized Level Replay (PLR) contribute to efficient learning but do not replace the need for automated high-quality feedback signals.
@@ 3.1 Sub-Trajectory Filtered Behavior Cloning (SFBC)
$\mathcal{D}_{SFBC}=\{(s_{t},a_{t})\mid\tau_{i}\in\mathcal{D},P_{VLM}(\tau_{i})\geq\alpha\text{ and }P_{VLM}(\tau_{i+1})\geq\alpha\}.$ (5)
@@
$\mathcal{L}_{SFBC}=-\mathbb{E}_{(s_{t},a_{t})\sim\mathcal{D}_{SFBC}}\left[P_{VLM}(\tau_{t})\log\pi_{\theta}(a_{t}|s_{t})\right].$ (6)
@@ 3.2 Comparison to Alternative SFO Methods
  • VLM+TD3+BC: Interprets $P_{VLM}(\tau)$ as a reward and applies TD3+BC (Fujimoto and Gu,, 2021), a competitive offline RL algorithm.
  • S-DPO: Applies Direct Preference Optimization (DPO) (Rafailov et al.,, 2023) to sub-trajectories using VLM-derived rankings.
@@ 4 Experiments
To evaluate the effectiveness of SFBC, we conduct experiments in the Pendulum-v1 environment, a standard continuous control benchmark. Our evaluation focuses on assessing SFBC’s ability to correctly identify and stitch together the most optimal segments from sub-optimal demonstrations while outperforming conventional offline RL baselines. While the VLM processes image-based observations to provide feedback, the learned policy itself conditions on lower-dimensional vector representations of the state, following standard RL formulations.
@@ 4.1 Dataset Construction
4.2 Baseline Comparisons
  • TD3+BC (GT): A strong offline RL baseline using TD3+BC (Fujimoto and Gu,, 2021), trained with access to ground-truth rewards.
  • AWAC (GT): Another standard offline RL algorithm, AWAC (Nair et al.,, 2020), trained on the dataset with ground-truth rewards.
Furthermore, SFBC outperforms VLM+TD3+BC, justifying our choice to treat $P_{VLM}$ as a weight rather than a reward. Conceptually, the weighting scheme can be interpreted as a value function $Q^{\pi_{\theta}}$, such that the weighted loss, $\mathcal{L}_{SFBC}$, implicitly optimizes a policy gradient. However, $\mathcal{L}_{SFBC}$ lacks an importance sampling correction, and $P_{VLM}$ is computed from offline data rather than the current policy, making it a better model of $Q^{\pi_{\text{behavior}}}$ (or $Q^{\pi_{\text{expert}}}$ after filtering). SFBC also resembles AWAC (Nair et al.,, 2020), except that the weighting $P_{VLM}$ should likewise be interpreted as approximating $e^{A^{\pi_{\text{expert}}}}$, rather than the advantage of the learned policy $e^{A^{\pi_{\theta}}}$. While SFBC does not fully correspond to a policy gradient, our results show that avoiding a direct reward interpretation—and the instability it introduces in reward propagation—is beneficial in practice.
@@ 5 Conclusion
Our findings suggest that VLM feedback provides non-myopic guidance, mitigating the need for explicit reward propagation and addressing key challenges in scaling offline RL. Most immediately, this work could be applied to automate feedback across a sufficiently diverse set of tasks to support training an RL foundation model from offline data. Such an application of SFBC would inherently fall under the category of offline meta-RL Beck et al., (2023), while also obviating the need for complex meta-gradient computation Vuorio et al., (2021). Putting such guidance systems on autopilot may finally provide the runway needed to arrive at a truly generalist agent for RL.

### v4

@@ Abstract
While internet-scale image and textual data have enabled strong generalization in Vision-Language Models (VLMs), the absence of internet-scale control data has impeded the development of similar generalization in standard reinforcement learning (RL) agents. Although VLMs are fundamentally limited in their ability to solve control tasks due to their lack of action-conditioned training data, their capacity for image understanding allows them to provide valuable feedback in RL tasks by recognizing successful outcomes. A key challenge in Reinforcement Learning from AI Feedback (RLAIF) is determining how best to integrate VLM-derived signals into the learning process. We explore this question in the context of offline RL and introduce a class of methods called Sub-Trajectory Filtered Optimization (SFO). We identify three key insights. First, trajectory length plays a crucial role in offline RL, as full-trajectory preference learning exacerbates the stitching problem, necessitating the use of sub-trajectories. Second, even in Markovian environments, a non-Markovian reward signal from a sequence of images is required to assess trajectory improvement, as VLMs do not interpret control actions and must rely on visual cues over time. Third, a simple yet effective approach—filtered and weighted behavior cloning—consistently outperforms more complex RLHF-based methods. We propose Sub-Trajectory Filtered Behavior Cloning (SFBC), a method that leverages VLM feedback on sub-trajectories while incorporating a retrospective filtering mechanism that removes sub-trajectories preceding failures to improve robustness.
@@ 1 Introduction
Reinforcement Learning from AI Feedback (RLAIF) presents a promising approach by leveraging vision-language models (VLMs) as scalable sources of learning signals (Bai et al., 2022; Klissarov et al., 2023; Ma et al., 2023; Faldor et al., 2024; Rocamonde et al., 2024; Baumli et al., 2023; Lee et al., 2024; Wang et al., 2024). However, determining how best to integrate VLM-derived signals into RL remains an open challenge. A key difficulty is that VLMs struggle to discriminate between similar trajectories, making it difficult to provide precise, informative feedback. This challenge is further complicated in online RL, where agents must learn from randomly initialized policies, leading to highly ambiguous early-stage feedback.
@@
    First, trajectory length is a crucial factor in offline RLAIF, since full-trajectory evaluation exacerbates the well-known stitching problem. This problem necessitates the use of sub-trajectories, as depicted in Figure 1.
@@
    Second, even in Markovian environments, a non-Markovian reward model based on image sequences is required to assess trajectory improvement. Since VLMs are not trained on action-conditioned data, they have no basis for reasoning about how actions influence state transitions and must instead rely on visual cues over time. This is depicted in Figure 1. While discrete actions can sometimes be represented textually, many control tasks involve continuous actions (e.g., torques) that cannot be meaningfully encoded in a format suitable for VLMs. Furthermore, providing a reward at every time step by querying a VLM is computationally intractable for many problems without additional distillation into a smaller model (Wang et al., 2024), and sub-sampling states to estimate reward would already abandon the Markov property.
@@
To that end, we introduce Sub-Trajectory Filtered Behavior Cloning (SFBC), which leverages VLM feedback on sub-trajectories while incorporating a retrospective filtering mechanism that removes sub-trajectories preceding failures. We empirically evaluate SFBC in a toy control domain, demonstrating its effectiveness compared to existing baselines, including naive behavior cloning and preference-based methods such as Direct Preference Optimization (DPO) (Rafailov et al., 2023). Notably, our method outperforms AWAC (Nair et al., 2020) and TD3+BC (Fujimoto & Gu, 2021), even when those methods use ground-truth rewards, highlighting the effectiveness of VLM-driven sub-trajectory feedback.
@@ 2 Related Work
One approach to improving RL efficiency is the manual collection of large-scale offline datasets (Vuong et al., 2023). However, this is difficult to scale due to the expense of generating diverse and high-quality trajectories on robotics in the real world. Similarly, human annotation of RL datasets is costly and difficult to scale, given the need for expert-level feedback and the challenge of obtaining consistent reward signals across tasks.
VLMs offer a promising alternative to human annotation by providing AI-generated feedback. Prior work has explored using VLMs for reward design and environment generation in code. Eureka and OMNI-EPIC (Ma et al., 2023; Faldor et al., 2024) have proposed methods where VLMs write programs to specify rewards and structure environments for learning. However, just as machine learning is needed to specify object categories (e.g., recognizing a dishwasher) rather than relying on hand-crafted rules, RL also requires machine-learned reward models rather than fixed rules from a VLM (e.g., to identify when a dishwasher has been unloaded). We firmly believe that hardcoded rules from a VLM cannot specify sufficiently complex tasks to elicit sufficiently robust reinforcement learning agents.
Recent work has also attempted to leverage AI-based reward detection to improve RLHF. Agent Q (Putta et al., 2024) uses AI-driven reward and success detection to generate preference labels for DPO. RL-VLM-F (Wang et al., 2024) distills a reward model from AI-generated labels but is an online method, and operates on Markovian states, which we find to be suboptimal. Rocamonde et al. (2024) and Baumli et al. (2023) use a CLIP model to provide Markovian feedback. Additionally, RoboCLIP (Sontakke et al., 2023) encodes text and video representations using CLIP to provide feedback. While RoboCLIP-style methods offer a useful way to encode feedback, we propose an end-to-end approach using a VLM (GPT-4o in our case), which provides greater flexibility. Our study focuses on how best to utilize this feedback and how to leverage feedback in an offline context.
Complementary approaches involve automatic environment generation and world modeling (Bruce et al., 2024; Valevski et al., 2024). These efforts provide a scalable way to generate diverse training environments, but they still require specifying reward signals to learn effectively. Similarly, automated task selection (Wang et al., 2023) and task ordering techniques such as Prioritized Level Replay (PLR) contribute to efficient learning but do not replace the need for automated high-quality feedback signals.
@@ 3.1 Sub-Trajectory Filtered Behavior Cloning (SFBC)
$\mathcal{D}_{SFBC}=\left\{(s_{t},a_{t},\tau_{i})\;\middle|\;\tau_{i}\in\mathcal{D},\;(s_{t},a_{t})\in\tau_{i},\;P{VLM}(\tau_{i})\geq\alpha,\;P_{VLM}(\tau_{i+1})\geq\alpha\right\}.$ (5)
@@
$\mathcal{L}_{SFBC}=-\mathbb{E}_{(s_{t},a_{t},\tau_{i})\sim\mathcal{D}_{SFBC}}\left[P_{VLM}(\tau_{i})\log\pi_{\theta}(a_{t}|s_{t})\right].$ (6)
@@ 3.2 Comparison to Alternative SFO Methods
  • VLM+TD3+BC: Interprets $P_{VLM}(\tau)$ as a reward and applies TD3+BC (Fujimoto & Gu, 2021), a competitive offline RL algorithm.
  • S-DPO: Applies Direct Preference Optimization (DPO) (Rafailov et al., 2023) to sub-trajectories using VLM-derived rankings.
@@ 4 Experiments
To evaluate the effectiveness of SFBC, we conduct experiments in the Pendulum-v1 environment, a standard continuous control benchmark. Our evaluation focuses on assessing SFBC’s ability to correctly identify and stitch together the most optimal segments from sub-optimal demonstrations while outperforming conventional offline RL baselines. While the VLM (ChatGPT-4o) processes image-based observations to provide feedback, the learned policy itself conditions on lower-dimensional vector representations of the state, following standard RL formulations.
@@ 4.1 Dataset Construction
4.2 Baseline Comparisons
  • TD3+BC (GT): A strong offline RL baseline using TD3+BC (Fujimoto & Gu, 2021), trained with access to ground-truth rewards.
  • AWAC (GT): Another standard offline RL algorithm, AWAC (Nair et al., 2020), trained on the dataset with ground-truth rewards.
Furthermore, SFBC outperforms VLM+TD3+BC, justifying our choice to treat $P_{VLM}$ as a weight rather than a reward. Conceptually, the weighting scheme can be interpreted as a value function $Q^{\pi_{\theta}}$, such that the weighted loss, $\mathcal{L}_{SFBC}$, implicitly optimizes a policy gradient. However, $\mathcal{L}_{SFBC}$ lacks an importance sampling correction, and $P_{VLM}$ is computed from offline data rather than the current policy, making it a better model of $Q^{\pi_{\text{behavior}}}$ (or $Q^{\pi_{\text{expert}}}$ after filtering). SFBC also resembles AWAC (Nair et al., 2020), except that the weighting $P_{VLM}$ should likewise be interpreted as approximating $e^{A^{\pi_{\text{expert}}}}$, rather than the advantage of the learned policy $e^{A^{\pi_{\theta}}}$. While SFBC does not fully correspond to a policy gradient, our results show that avoiding a direct reward interpretation—and the instability it introduces in reward propagation—is beneficial in practice.
@@ 5 Conclusion
Our findings suggest that VLM feedback provides non-myopic guidance, mitigating the need for explicit reward propagation and addressing key challenges in scaling offline RL. Most immediately, this work could be applied to automate feedback across a sufficiently diverse set of tasks to support training an RL foundation model from offline data. Such an application of SFBC would inherently fall under the category of offline meta-RL (Beck et al., 2023), while also obviating the need for complex meta-gradient computation (Vuorio et al., 2021). Putting such guidance systems on autopilot may finally provide the runway needed to arrive at a truly generalist agent for RL.
@@
Supplementary Materials
The following content was not necessarily subject to peer review.